# Emotion-LLaMA: Multimodal Emotion Recognition and Reasoning with Instruction Tuning

### v1

@@ Abstract
Accurate emotion perception is crucial for various applications, including human-computer interaction, education, and counseling. However, traditional single-modality approaches often fail to capture the complexity of real-world emotional expressions, which are inherently multimodal. Moreover, existing Multimodal Large Language Models (MLLMs) face challenges in integrating audio and recognizing subtle facial micro-expressions. To address this, we introduce the MERR dataset, containing 28,618 coarse-grained and 4,487 fine-grained annotated samples across diverse emotional categories. This dataset enables models to learn from varied scenarios and generalize to real-world applications. Furthermore, we propose Emotion-LLaMA, a model that seamlessly integrates audio, visual, and textual inputs through emotion-specific encoders. By aligning features into a shared space and employing a modified LLaMA model with instruction tuning, Emotion-LLaMA significantly enhances both emotional recognition and reasoning capabilities. Extensive evaluations show Emotion-LLaMA outperforms other MLLMs, achieving top scores in Clue Overlap (7.83) and Label Overlap (6.25) on EMER, an F1 score of 0.9036 on MER2023 challenge, and the highest UAR (45.59) and WAR (59.37) in zero-shot evaluations on DFEW dataset.
@@ 1 Introduction
Emotion perception plays a vital role in applications such as human-computer interaction [13, 14, 15, 58, 51], educational assistance [31], and psychological counseling [6, 30]. While single-modality approaches, including facial expression recognition [33, 64, 52], text emotion analysis [37, 29, 20], and audio emotion recognition [23, 35, 27], have shown effectiveness, real-world emotional data is often multimodal, integrating text, audio, and images.
Despite extensive multimodal fusion methods having achieved promising improvements [10, 11, 12, 41, 42, 43, 68, 72, 75, 78], they mainly focus on feature interaction and modality completion, remaining under-explored for knowledge-level interaction which is essential for emotional reasoning of humans. Recently, Multimodal Large Language Models (MLLMs) have excelled in tasks such as visual-language understanding [48], visual question answering [79], and video understanding [46, 63]. However, for emotion recognition [33], models like GPT-4 with Vision (GPT-4V) still face two main challenges: the inability to process audio and the failure to recognize micro-expressions.
@@
To address these challenges, we introduce the MERR dataset (Sec. 3.1), which enables multimodal large models and supports instruction tuning to learn from diverse scenarios and generalize to real-world applications. We also propose the Emotion-LLaMA model (Sec. 3.2), which integrates audio, visual, and textual inputs through emotion-specific encoders. By employing instruction tuning (Sec. 3.3), Emotion-LLaMA significantly enhances both the accuracy of emotional recognition and the depth of emotional reasoning, setting a new benchmark for multimodal emotion analysis. Extensive experiments and evaluations (Sec. 4) demonstrate Emotion-LLaMA’s superiority, achieving top scores on EMER, MER2023111http://merchallenge.cn/mer2023, and DFEW datasets. Our main contributions are as follows:
@@
  • We developed the Emotion-LLaMA model, which incorporates HuBERT for audio processing and multiview visual encoders (MAE, VideoMAE, EVA) for capturing facial details, dynamics, and context. By aligning these features into a modified LLaMA language model, Emotion-LLaMA enhances emotional recognition and reasoning capabilities. The model excels in multimodal emotion recognition and reasoning, particularly through the innovative use of instruction tuning, which significantly improves its performance.
  • Extensive experiments demonstrate that Emotion-LLaMA significantly outperforms other MLLMs across multiple datasets, establishing it as the current state-of-the-art model in public competitions. It achieved the highest Clue Overlap score of 7.83 and Label Overlap score of 6.25 on the EMER dataset. Additionally, Emotion-LLaMA attained an impressive F1 score of 0.9036 on the MER2023 dataset. Furthermore, in zero-shot evaluations on the DFEW dataset, it achieved the highest Unweighted Average Recall (UAR) of 45.59 and Weighted Average Recall (WAR) of 59.37.
@@ 3.1 MERR Dataset Construction
The MERR dataset is constructed through a comprehensive process of emotional expression annotation in video data, as outlined in Algorithm 1 and Figure 1. First, facial features are extracted from each video frame using the OpenFace toolkit, which detects and scores Action Units (AUs) to identify the frame with the maximum cumulative intensity:
@@
where $S_{au_{i}}$ represents the intensity of each AU. These AUs are mapped to facial expression descriptions $C_{ved}$ (Tables 7 and 8) to accurately depict facial movements. Next, MiniGPT-v2222https://github.com/Vision-CAIR/MiniGPT-4/blob/main/demo_v2.py analyzes the peak frame to extract contextual information $C_{vod}$, such as activities and environment (Figure 1), facilitating the identification of latent emotional elements within the background context. Qwen-Audio [^3] https://www.modelscope.cn/models/qwen/QWen-Audio/summary processes audio segments to extract nuances in speech and vocal tone, generating emotion-related descriptions $C_{atd}$. Visual and audio information are concatenated into a raw multimodal description, integrating sensory inputs to enhance the contextual supplementation for lexical subtitles. Lexical subtitles $C_{ls}$ are integrated into the multimodal description, providing textual context that complements the audio and visual data. LLaMA-3444https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct refines these annotations by aggregating unimodal descriptions ($C_{ved}$, $C_{vod}$, $C_{atd}$, $C_{ls}$) into a detailed multimodal description $C_{md}$, following instructions and examples in Table 9. Finally, the comprehensive description $C_{md}$ is used to annotate the peak frame, ensuring the video is annotated with detailed emotional descriptors.
@@
17:Annotate $Frame_{\text{peak}}$ with $C_{md}$ return Annotated $Frame_{\text{peak}}$
The MERR dataset extends the range of emotional categories and annotations beyond those found in existing datasets (Table 12). Each sample is annotated with an emotion label and described in terms of its emotional expression. The dataset was initially auto-annotated with coarse-grained labels for 28,618 samples from a large pool of unannotated data, and later refined to include 4,487 samples with fine-grained annotations. Figure 5 shows that, compared to other datasets, MERR encompasses a wider range of emotional categories, including easily confusable ones like “doubt” and “contempt”. More details of MERR dataset construction are provided in the project homepage.
@@ 3.2 Multimodal Emotion-LLaMA Model
Multimodal Prompt Template. To address the intricate needs of emotional understanding, we craft a structured multimodal prompt template incorporating descriptive captions and emotion flags (as detailed in Table 13 and  14), directing the LLM to decipher latent correlations between emotional states and corresponding visual or auditory content. The template is denoted as:
[INST] $<$ AudioFeature$>$ $<$VideoFeature$>$ [Task Identifier] Prompt [/INST]
Multiview Multimodal Encoder. To capture emotional cues in audio and visual modalities, we leverage the HuBERT  [27] model as our audio encoder $\mathcal{E}^{aud}$ and a multiview visual encoder $\mathcal{E}^{vis}$. HuBERT extracts a comprehensive auditory representation $u^{aud}$ from the input audio signal $A$, exhibiting remarkable performance in emotion recognition tasks.
@@ 3.3 Training of Emotion-LLaMA Model
We design a multi-task learning scheme to simultaneously supervise the model in learning emotional reasoning and recognition. The ground truth output and labels are converted and concatenated as standard text by a formatted template for autoregressive loss calculation  [74]. Iterative random instruction sampling (see Table 13 and 14 for full list of instructions) for emotional reasoning and recognition tasks during training guides the model to learn a comprehensive understanding of emotional reasoning. Typically, Emotion-LLaMA is trained in a coarse-to-fine manner, consisting of the Pre-training and Multimodal Instruction Tuning:
Stage 1: Pretraining. Initially, the model is trained on coarse-grained data, including visual and audio descriptions. Distinct task identifiers help the model grasp emotions from multiple perspectives. This phase involves simple descriptions or classifications, facilitating the rapid alignment of multimodal feature tokens ($\langle\text{T}^{aud}\rangle$, $\langle\text{T}^{vis}_{glo}\rangle$, $\langle\text{T}^{vis}_{loc}\rangle$, and $\langle\text{T}^{vis}_{temp}\rangle$) to the word embedding space  [8, 73]. More details are in the code repository.
Stage 2: Multimodal Instruction Tuning. The pretrained Emotion-LLaMA model is then refined using fine-grained instructional datasets to enhance its capacity for emotion recognition and reasoning. This stage utilizes multimodal instruction tuning datasets, incorporating comprehensive reasoning descriptions from the MERR dataset. The tuning process is extended to diverse sources, including MER2023  [43] and DFEW  [33], which feature precisely annotated emotional categories. This phase ensures that the model not only identifies emotions accurately but also understands the underlying context and reasoning behind each emotion. More details are in the code repository.
@@ 4.1 Experimental Setup
To verify the effectiveness of Emotion-LLaMA, we conducted extensive evaluations across four different datasets: MERR, MER2023  [43], DFEW  [33], and EMER  [45]. Notably, we utilized the MERR dataset for pre-training the model and then fine-tuned it on target datasets for evaluation.
Emotion Recognition Evaluation. We performed instruction-tuning on the MER2023  [43] and DFEW  [33] datasets, allowing the model to integrate the emotional knowledge acquired during pre-training for enhanced emotion reasoning. To test the generalizability of our model, we used three datasets: MER2023, DFEW, and EMER. Both MER2023 and DFEW are multimodal emotion recognition datasets composed of movie and TV series clips, each annotated with various emotion categories. For fair comparisons, we evaluated Emotion-LLaMA on MER2023-SEMI using the F1 score and compared it with other MLLMs and state-of-the-art (SOTA) methods using unweighted average recall (UAR) and weighted average recall (WAR).
@@ 4.3 Comparison with State-of-the-Art Methods
Multimodal Emotion Reasoning Results. We compared Emotion-LLaMA with contemporary MLLMs such as Video-LLaMA  [73], Video-ChatGPT  [50], PandaGPT  [59], VideoChat  [40], and Valley  [49], presenting the results in Table 4.3. VideoChat demonstrates that aligning visual data directly with textual embedding space (VideoChat-Embed) significantly outperforms converting it into textual format (VideoChat-Text), supporting our method of mapping audio and visual features to textual embedding space. Notably, other MLLMs that accept audio inputs, like PandaGPT and Video-LLaMA, show no standout performance, suggesting inefficiencies in extracting rich emotional content from audio. Emotion-LLaMA excels beyond these models across both Clue Overlap and Label Overlap evaluation metrics, highlighting our model’s unparalleled ability to extract direct emotional features and engage in logical emotional reasoning. The scoring criteria and cases are presented in the appendix.
Multimodal Emotion Recognition Results. Table 2 presents the comparison results on the DFEW dataset. In the zero-shot scenario, Emotion-LLaMA demonstrates superior capabilities compared to all other MLLMs, showcasing its strong generalization ability. After fine-tuning, Emotion-LLaMA achieves the highest Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) scores, further indicating its exceptional performance in emotion recognition tasks. These results highlight the effectiveness of our model in adapting to new datasets and accurately identifying emotions across various modalities. Overall, the results on multiple datasets of Emotion-LLaMA’s performance highlight the effectiveness of our approach in accurately recognizing emotions from multimodal data.
@@ 4.4 Multimodal Emotion Recognition Challenge
To further validate the effectiveness of our proposed Emotion-LLaMA model, we conducted experiments on the MER2023 Challenge [^5] http://merchallenge.cn/mer2023 dataset, comparing it with previous state-of-the-art supervised methods. The results, presented in Table 4.4, demonstrate that our model, which maps audio and visual features to the textual space, achieves the highest F1 score across various modalities.
Table 3: Comparison of multimodal emotion recognition results on MER2023.
@@
Emotion-LLaMA (ours) A, V 0.8905
Emotion-LLaMA (ours) A, V, T 0.9036
This approach significantly enhances the context of the textual modality by providing a more comprehensive understanding of the information, thereby outperforming other models. By integrating audio, visual, and textual data, Emotion-LLaMA can better capture the nuances of emotional expression, leading to more accurate and reliable emotion recognition. These results showcase the robustness and versatility of our approach in handling complex multimodal data for emotion recognition tasks, making it a promising solution for real-world applications.

### v2

@@ Abstract
Accurate emotion perception is crucial for various applications, including human-computer interaction, education, and counseling. However, traditional single-modality approaches often fail to capture the complexity of real-world emotional expressions, which are inherently multimodal. Moreover, existing Multimodal Large Language Models (MLLMs) face challenges in integrating audio and recognizing subtle facial micro-expressions. To address this, we introduce the MERR dataset, containing 28,618 coarse-grained and 4,487 fine-grained annotated samples across diverse emotional categories. This dataset enables models to learn from varied scenarios and generalize to real-world applications. Furthermore, we propose Emotion-LLaMA, a model that seamlessly integrates audio, visual, and textual inputs through emotion-specific encoders. By aligning features into a shared space and employing a modified LLaMA model with instruction tuning, Emotion-LLaMA significantly enhances both emotional recognition and reasoning capabilities. Extensive evaluations show Emotion-LLaMA outperforms other MLLMs, achieving top scores in Clue Overlap (7.83) and Label Overlap (6.25) on EMER, an F1 score of 0.9036 on MER2023-SEMI challenge, and the highest UAR (45.59) and WAR (59.37) in zero-shot evaluations on DFEW dataset.
@@ 1 Introduction
Emotion perception plays a vital role in applications such as human-computer interaction [17, 18, 19, 80], educational assistance [43], and psychological counseling [7, 42]. While single-modality approaches, including facial expression recognition [45, 57, 70, 87], text emotion analysis [25, 49, 41], and audio emotion recognition [31, 39, 47], have shown effectiveness, real-world emotional data is often multimodal, integrating text, audio, and images.
Despite extensive multimodal fusion methods having achieved promising improvements [13, 14, 16, 56, 58, 59, 77, 91, 96, 100, 103], they mainly focus on feature interaction and modality completion, remaining under-explored for knowledge-level interaction which is essential for emotional reasoning of humans. Recently, Multimodal Large Language Models (MLLMs) have excelled in tasks such as visual-language understanding [50, 65], visual question answering [104], and video understanding [35, 63, 86]. However, for emotion recognition [45], models like GPT-4 with Vision (GPT-4V) still face two main challenges: the inability to process audio and the failure to recognize micro-expressions.
@@
To address these challenges, we introduce the MERR dataset (Sec. 3.1), which enables multimodal large models and supports instruction tuning to learn from diverse scenarios and generalize to real-world applications. We also propose the Emotion-LLaMA model (Sec. 3.2), which integrates audio, visual, and textual inputs through emotion-specific encoders. By employing instruction tuning (Sec. 3.3), Emotion-LLaMA significantly enhances both the accuracy of emotional recognition and the depth of emotional reasoning, setting a new benchmark for multimodal emotion analysis. Extensive experiments and evaluations (Sec. 4) demonstrate Emotion-LLaMA’s superiority, achieving top scores on EMER, MER2023111http://merchallenge.cn/mer2023, MER2024222https://zeroqiaoba.github.io/MER2024-website, and DFEW datasets. Our main contributions are as follows:
@@
  • We developed the Emotion-LLaMA model, which incorporates HuBERT for audio processing and multiview visual encoders (MAE, VideoMAE, EVA) for capturing facial details, dynamics, and context. By aligning these features into a modified LLaMA language model, Emotion-LLaMA enhances emotional recognition and reasoning capabilities.
  • Extensive experiments demonstrate that Emotion-LLaMA significantly outperforms other MLLMs across multiple datasets, establishing it as the current state-of-the-art model in public competitions. It achieved top scores on the EMER dataset (Clue Overlap: 7.83, Label Overlap: 6.25) and attained F1 scores of 0.9036 on MER2023-SEMI11footnotemark: 1 and 0.8452 on MER2024-NOISE22footnotemark: 2. Emotion-LLaMA also surpassed ChatGPT-4V in zero-shot evaluations, including DFEW (+4.37%) and MER2024-OV22footnotemark: 2 (+8.52%).
@@ 3.1 MERR Dataset Construction
The MERR dataset is constructed through a comprehensive process of emotion annotation in video data, as outlined in Algorithm 1 and Figure 1. First, human faces are extracted from each video frame using the OpenFace toolkit, which detects and scores Action Units (AUs) [28, 69] to identify the frame with the maximum cumulative intensity:
@@
where $S_{au_{i}}$ represents the intensity of each AU. These AUs are mapped to facial expression descriptions $C_{ved}$ (Tables 10 and 11) to accurately depict facial movements. Next, MiniGPT-v2333https://github.com/Vision-CAIR/MiniGPT-4/blob/main/demo_v2.py analyzes the peak frame to extract contextual information $C_{vod}$, such as activities and environment (Figure 1), facilitating the identification of latent emotional elements within the background context. Qwen-Audio [^4] https://www.modelscope.cn/models/qwen/QWen-Audio/summary processes audio segments to extract nuances in speech and vocal tone, generating emotion-related descriptions $C_{atd}$. Visual and audio information are concatenated into a raw multimodal description, integrating sensory inputs to enhance the contextual supplementation for lexical subtitles. Lexical subtitles $C_{ls}$ are integrated into the multimodal description, providing textual context that complements the audio and visual data. LLaMA-3555https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct refines these annotations by aggregating unimodal descriptions ($C_{ved}$, $C_{vod}$, $C_{atd}$, $C_{ls}$) into a detailed multimodal description $C_{md}$, following instructions and examples in Table 12. Finally, the comprehensive description $C_{md}$ is used to annotate the peak frame, ensuring the video is annotated with detailed emotional descriptors.
@@
17:return $C_{md}$
The MERR dataset extends the range of emotional categories and annotations beyond those found in existing datasets (Table 15). Each sample is annotated with an emotion label and described in terms of its emotional expression. The dataset was initially auto-annotated with coarse-grained labels for 28,618 samples from a large pool of unannotated data using LLaMA-3, and was later refined to include 4,487 samples with fine-grained annotations, carefully selected by experts. Figure 5 shows that, compared to other datasets, MERR encompasses a wider range of emotional categories. More details of MERR dataset construction are provided in the project homepage and the Appendix A.
@@ 3.2 Multimodal Emotion-LLaMA Model
Multimodal Prompt Template. To address the intricate needs of emotional understanding, we craft a structured multimodal prompt template incorporating descriptive captions and emotion flags (as detailed in Table 16 and  17), directing the LLM to decipher latent correlations between emotional states and corresponding visual or auditory content. The template is denoted as:
[INST] $<$VideoFeature$>$ $<$AudioFeature$>$ [Task Identifier] Prompt [/INST]
Multiview Multimodal Encoder. To capture emotional cues in audio and visual modalities, we leverage the HuBERT  [39] model as our audio encoder $\mathcal{E}^{aud}$ and a multiview visual encoder $\mathcal{E}^{vis}$. HuBERT extracts a comprehensive auditory representation $u^{aud}$ from the input audio signal, exhibiting remarkable performance in emotion recognition tasks.
@@ 3.3 Training of Emotion-LLaMA Model
We design a multi-task learning scheme to simultaneously supervise the model in learning emotional reasoning and recognition. The ground truth output and labels are converted and concatenated as standard text by a formatted template for autoregressive loss calculation  [99]. Iterative random instruction sampling (see Table 16 and 17 for full list of instructions) for emotional reasoning and recognition tasks during training guides the model to develop a comprehensive understanding of emotions. Typically, Emotion-LLaMA is trained in a coarse-to-fine manner, consisting of the Pre-training and Multimodal Instruction Tuning:
Stage 1: Pretraining. Initially, the model is trained on 28,618 coarse-grained samples from the MERR dataset. Distinct tasks help the model grasp emotions from multiple perspectives. This phase involves simple descriptions or classifications, facilitating the rapid alignment of multimodal feature tokens ($\langle\text{T}^{aud}\rangle$, $\langle\text{T}^{vis}_{glo}\rangle$, $\langle\text{T}^{vis}_{loc}\rangle$, and $\langle\text{T}^{vis}_{temp}\rangle$) to the word embedding space  [10, 97].
Stage 2: Multimodal Instruction Tuning. The pretrained Emotion-LLaMA model is then refined using fine-grained instructional datasets to enhance its capacity for emotion recognition and reasoning. This stage utilizes multimodal instruction tuning datasets, incorporating 4,487 fine-grained annotated descriptions for comprehensive reasoning from the MERR dataset. The tuning process is extended to diverse sources, including MER2023  [59] and DFEW  [45], which feature precisely annotated emotional categories. This phase ensures that the model not only identifies emotions accurately but also understands the underlying context and reasoning behind each emotion. More details are in the code repository and the Appendix B.
@@ 4.1 Experimental Setup
To verify the effectiveness of Emotion-LLaMA, we conducted extensive evaluations across four different datasets: MER2023  [59], MER2024  [60], DFEW  [45], and EMER  [62]. Notably, we utilized the MERR dataset for pre-training the model and then fine-tuned it on target datasets for evaluation.
Emotion Recognition Evaluation. We performed instruction tuning on the MER2023 and DFEW datasets, allowing the model to integrate the emotional knowledge acquired during pretraining. To test the generalizability of our model, we used three datasets: MER2023, MER2024, and DFEW. These datasets are multimodal emotion recognition datasets composed of movie and TV series clips, each annotated with various emotion categories. For fair comparisons, we evaluated Emotion-LLaMA on MER2023-SEMI and MER2024-NOISE using the F1 score. We also compared it with other MLLMs and state-of-the-art (SOTA) methods using unweighted average recall (UAR) and weighted average recall (WAR) on the DFEW dataset. Additionally, we used the average of accuracy and recall scores as evaluation metrics on the MER2024-OV dataset.
@@ 4.3 Comparison with State-of-the-Art Methods
Multimodal Emotion Reasoning Results. We compared Emotion-LLaMA with contemporary MLLMs such as Video-LLaMA, Video-ChatGPT, PandaGPT, VideoChat, and Valley, presenting the results in Table 4.3. VideoChat demonstrates that aligning visual data directly with textual embedding space (VideoChat-Embed) significantly outperforms converting it into textual format (VideoChat-Text), supporting our method of mapping audio and visual features to textual embedding space. Notably, other MLLMs that accept audio inputs, like PandaGPT and Video-LLaMA, show no standout performance, suggesting inefficiencies in extracting rich emotional content from audio. Emotion-LLaMA excels beyond these models across both Clue Overlap and Label Overlap evaluation metrics, highlighting our model’s unparalleled ability to extract direct emotional features and engage in logical emotional reasoning. The scoring criteria and cases are presented in the Appendix C.1.
Multimodal Emotion Recognition Results. Table 2 presents the comparison results on the DFEW dataset. In the zero-shot scenario, Emotion-LLaMA demonstrates superior capabilities compared to all other MLLMs, showcasing its strong generalization ability. Notably, the majority of MLLMs scored zero in the disgust category, with GPT-4V achieving only 10.34%. This may be attributed to safety constraints on the term "disgust" within large language models, indicating a need for further exploration. Additionally, different MLLMs tend to favor predicting a specific emotion category, resulting in higher scores for those categories but lower recall scores overall. In contrast, Emotion-LLaMA maintains a more balanced prediction across all categories, ultimately achieving the highest WAR score of 59.37%. After fine-tuning, Emotion-LLaMA achieves the highest Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) scores, further indicating its exceptional performance in emotion recognition tasks. These results highlight the effectiveness of our model in adapting to new datasets and accurately identifying emotions across various modalities. Overall, the results of Emotion-LLaMA’s performance highlight the effectiveness of our approach in accurately recognizing emotions from multimodal data.
@@ 4.4 Multimodal Emotion Recognition Challenge
To further validate the effectiveness of our proposed Emotion-LLaMA model, we conducted experiments on the MER2023 and MER2024 Challenge, comparing it with previous state-of-the-art methods.The results, presented in Table 4.4, demonstrate that our model, which maps audio and visual features to the textual space, achieves the highest F1 score across various modalities. This approach significantly enhances the context of the textual modality by providing a more comprehensive understanding of the information, thereby outperforming other models. By integrating audio, visual, and textual data, Emotion-LLaMA can better capture the nuances of emotional expression, leading to more accurate and reliable emotion recognition.
The MER2024 Challenge introduced a new Open-Vocabulary Multimodal Emotion Recognition (MER-OV) task. Unlike traditional tasks, MER-OV focuses on recognizing any number of labels across diverse categories, aiming for a more nuanced and precise description of emotional states. As shown in Table 4.4, Emotion-LLaMA outperforms other mainstream multimodal models, yielding an 8.52% improvement in average accuracy and recall compared to GPT-4V, and achieving the highest zero-shot score among all participating large multimodal models. These results showcase the robustness and versatility of our approach in handling complex multimodal data for emotion recognition tasks, making it a promising solution for real-world applications.
Table 3: Comparison of multimodal emotion recognition results on MER2023. The table shows the performance of different models across various modalities, with the highest F1 scores achieved by our proposed method.
@@
wav2vec 2.0  [4] A 0.4028
VGGish  [38] A 0.5481
HuBERT  [39] A 0.8511
ResNet  [37] V 0.4132
MAE  [36] V 0.5547
VideoMAE  [84] V 0.6068
RoBERTa  [66] T 0.4061
BERT  [25] T 0.4360
MacBERT  [24] T 0.4632
@@
Emotion-LLaMA A, V 0.8905
Emotion-LLaMA A, V, T 0.9036
Table 4: Performance (%) of Multimodal Large Language Models on MER2024 Challenge track 3: MER-OV. The “avg” column represents the average of “$\mbox{Accuracy}_{\mbox{s}}$” and “$\mbox{Recall}_{\mbox{s}}$
Model $\mbox{Accuracy}_{\mbox{s}}$ $\mbox{Recall}_{\mbox{s}}$ Avg
Empty 0.00 0.00 0.00
Random 13.42 24.85 19.13
Ground Truth 93.37 52.51 72.94
Valley  [67] 20.16 13.26 16.71
Otter  [51] 29.64 23.04 26.34
PandaGPT  [81] 35.75 31.57 33.66
Video-LLaMA  [97] 31.08 32.26 31.67
VideoChat  [54] 43.17 44.92 44.05
VideoChat2  [55] 46.91 34.78 40.85
Video-ChatGPT  [68] 46.20 39.33 42.77
SALMONN  [83] 42.20 44.75 43.47
Qwen-Audio  [22] 55.12 32.91 44.02
mPLUG-Owl  [94] 44.80 46.54 45.67
AffectGPT  [62] 66.14 46.56 56.35
GPT-4V  [71] 56.19 58.97 57.58
Emotion-LLaMA 69.61 62.59 66.10